# Agentic Structured Graph Traversal for Root Cause Analysis of Code-related Incidents in Cloud Applications

### v1

@@ VI-B Ablation Study (RQ2)
    This variant disables PDG-traversal and program context ($C_{P}$), so only the incident context and observability context ($c_{i}$) are used in the RCA decision in §IV-C. This experiment characterized the performance gains resulting from the use of program context.
@@
    The Raw Code variant skips PDG construction (§IV-A) and LLM-driven PDG traversal (§IV-C) when constructing $C_{P}$ and instead uses an LLM with a modified $\psi_{syn}$ prompt (§IV-C) to consolidate and generate program context directly from aggregated raw code files. We limited the raw code to 800k characters to prevent context overflow.
Evaluation setup. We evaluated these Praxis variants using the best-performing LLM for Praxis–gpt-5-codex. For statistical robustness, we repeated each $\langle\text{variant},\ \text{scenario}\rangle$ experiment five times with distinct random seeds, yielding an additional 300 trajectories.
Figure 7: RCA reasoning of Praxis (Obs. Ctx.) and Praxis. Nodes are microservices; solid arrows are dependencies; the green dotted arrow is a dependency with missing traces and had to be derived from program context.
@@
Why do program context and LLM-driven PDG help? Similar to the SRE-Agent baseline, Praxis without program context (Praxis (Obs. Ctx.)) failed to identify root causes within, or hinted by, the internal code logic, resulting in superficial symptom-level RCA. Although the Praxis variant without PDG (Praxis (Raw Code)) retains service-level dependency awareness, the absence of PDG-guided diagnosis halved its accuracy compared to that of the full approach. Lacking explicit awareness of program-level dependencies, Praxis (Raw Code) cannot effectively focus its reasoning on code regions relevant to the incident. Instead, it “fixated” on ostensibly faulty regions–code that appears erroneous on the surface (e.g., manually thrown exceptions)–that are not pertinent to the observed failure. With code context exceeding 200k tokens at times, Praxis (Raw Code) exhibited needle in a haystack [gola_multi_needle_2024] and context rot [hong2025contextrot] phenomena, leading to degraded RCA accuracy. These findings confirm that PDG-guided diagnosis is essential: it enforces the alignment of observability data with concise, accurate code context structurally presented to the LLM, grounding the RCA in relevant code paths and code regions to yield the observed performance gains.
@@ VI-C Diagnosis Overheads (RQ3)
For this RQ, we evaluated the RCA overheads of SRE-Agent baselines, Praxis, Praxis (Obs. Ctx.), and Praxis (Raw Code) across all scenarios and using the best-performing LLMs: gpt-5 for SRE agent baselines, and gpt-5-codex for Praxis.
Evaluation setup. To account for differences in RCA reasoning accuracy, we normalized both MTTC and ATC by RCR Pass@1% to assess diagnosis efficiency: $\text{{Normalized metric}}=\frac{\text{{MTTC or ATC}}}{\text{RCR Pass@1\%}}$. The resulting normalized MTTC (i.e., mean time to diagnosis or MTTD) and normalized ATC (effective ATC) reflect time and token costs per successful diagnosis.
Table˜V shows that, when accounting for RCA accuracy, Praxis reduces MTTD by up to 24.8% [^7] Includes LLM backend response times. relative to SRE-Agent baselines. Moreover, compared to Praxis, the SRE baselines consume up to $3.8\times$ as many tokens per successful diagnosis, indicating substantially lower token efficiency.
When we compare Praxis with its ablated variants (Praxis (Raw Code)) and Praxis (Obs. Ctx.)), construction of program context via LLM-driven PDG traversal reduced MTTD and effective ATC by up to $85.9\%$ and $83.5\%$, respectively, highlighting their significant contributions to efficient RCA. Notably, LLM-driven PDG traversal resulted in a $6.1\times$ reduction in ATC compared to that of supplying raw code as plain text in prompt (as in [rcacp_chen2024, li2025coca_icse]). Praxis’s MTTD of 1,474 seconds is substantially lower than that of manual RCA, which can take several hours for incidents with comparable root causes.
@@
  • Metrics: (i) MTTC = mean time to completion in seconds; (ii) ATC = average token consumptions in tokens; (iii) MTTD = mean time to diagnosis; (iv) Eff. ATC = effective ATC. Best-performing variant is shown in bold; second-best is underlined. Lower is better.
VI-D Limitations and Future Work
@@
Scenarios with weak observability. Praxis’s effectiveness is contingent on the availability and quality of observability data to anchor its PDG traversal. Sparse traces, logs, metrics, or events can limit Praxis’s ability to accurately align runtime error signals with corresponding dependency paths to traverse. This limitation is most apparent for Deployment Manifest Error and Resource-related Fault incidents, where Praxis achieves near-zero accuracy (omitted from Table˜IV). Praxis tends to underperform in these cases because the fault often lies in the deployment manifest or third-party services, offering no direct observable signal to anchor RCA decision-making. Consequently, Praxis tends to falsely attribute the incident to healthy code paths, leading to false positives. While collecting exhaustive observability data could mitigate this, it would significantly incur production overheads.
Graph representation. Praxis’s effectiveness relies on the correctness and completeness of the dependency graphs. Dependency graphs that are incorrect or incomplete (e.g., an outdated PDG, or a missing microservice in SDG) hinder Praxis’s ability to localize the root cause. Moreover, Praxis is likely sensitive to the granularity of the node. For example, including large code context (module-level hammock block) can reduce diagnostic specificity, while a fine-grained representation (statement-level hammock block) risks missing broader context beneficial to RCA.
@@
Code availability and static code analysis. Praxis’s gains from PDG traversal are naturally limited if microservice code is unavailable for graph construction and it must rely solely on observability data for diagnosis. Conversely, conditioning too strongly on static code context risks unintended bias, where Praxis may mistakenly attribute faults to code blocks when the incident actually stems from external factors (e.g., deployment or resources). This risk can be mitigated by incorporating runtime instrumentation to confirm whether implicated code paths were indeed executed at the time of the incident.
Future Work. Future iterations of Praxis will be extended to platform- and infrastructure-level incidents. We also plan to integrate dynamic program analysis–incorporating runtime tracing, heap snapshots, and dynamic slicing–to ground static analysis in verifiable runtime behavior, enabling more comprehensive and precise RCA.
@@ VII Related Work
Agentic approach for cloud incident RCA. Recent agentic RCA approaches [roy2024exploring_rca_agents, jha2025itbench, xu2025openrca, chen2025aiopslab, chen2025stratus] adopted the ReAct paradigm [yao2023react] to analyze observability data but lack structured reasoning workflow and program context, limiting their effectiveness on software and misconfiguration faults. The prior approaches that include program/code context [wang2024rcagent, li2025coca_icse] inline code in plain-text prompts without explicitly imposing the inherent code structure, as they do not employ a PDG-guided reasoning process. Praxis advances beyond those efforts by performing LLM-driven reasoning over service dependency and program dependence graphs, unifying incident, observability, cloud, and program context for scalable, end-to-end RCA.
@@ VIII Conclusion
This paper proposes Praxis, an agentic approach that reasons over microservice and program dependency graphs for a comprehensive, in-depth root-cause analysis. Our evaluation of Praxis on 30 scenarios that span software, configuration, deployment, and resource failures demonstrates that Praxis achieves up to $3.1\times$ better root-cause reasoning accuracy, $2.1\times$ higher root-cause identification accuracy, and a $3.8\times$ reduction in token consumption compared to state-of-the-art agentic baselines.
@@ Acknowledgment
We thank R. Arora, B. Bhavya, N. Zheutlin, P. T. Isaza, J. Ahn, A. Paradkar, L. Shwartz, R. K. Kottapalli, A. D. Angelis, A. Patke, P. Cao, Z. Zheng, and J. Applequist for technical input and feedback on early drafts of this work, and H. C. Fairow, S. Weick, K. Atchley, Y. Deng, R. Pavuluri, M. Vukovic, X. Liu, D. Sow, and N. Fuller for administrative support. This work is supported by the IBM-Illinois Discovery Accelerator Institute (IIDAI). Any opinions expressed here are those of the authors and do not necessarily reflect the views of IBM.

### v2

@@ VI-B Ablation Study (RQ2)
    The observability context-only variant disables PDG-traversal and program context ($C_{P}$), so only the incident context and observability context ($c_{i}$) are used in the RCA decision in §IV-C. This experiment characterized the performance gains resulting from the use of program context.
@@
    The Raw Code variant skips PDG construction (§IV-A) and LLM-driven PDG traversal (§IV-C) when constructing $C_{P}$ and instead uses an LLM with a modified $\psi_{syn}$ prompt (§IV-C) to consolidate and generate program context directly from aggregated raw code files. We limited the raw code to 800k characters ($\sim$200 tokens) to avoid overflowing the LLM’s context window.
Evaluation setup. We evaluated these Praxis variants using the best-performing LLM for Praxis–gpt-5-codex. For statistical robustness, we repeated each $\langle\text{variant},\ \text{scenario}\rangle$ experiment five times with distinct random seeds, yielding a total of 300 trajectories.
Figure 7: RCA reasoning of Praxis (Obs. Ctx.) and Praxis. Nodes are microservices; solid arrows are dependencies; the green dotted arrow is a dependency with missing traces and had to be derived from program context. SVC. = Service.
@@
Why do program context and LLM-driven PDG help? Similar to the SRE-Agent baseline, Praxis without program context (Praxis (Obs. Ctx.)) failed to identify root causes within, or hinted by, the internal code logic, resulting in superficial symptom-level RCA. Although the Praxis variant without PDG (Praxis (Raw Code)) retains service-level dependency awareness, the absence of PDG-guided diagnosis halved its accuracy compared to that of the full approach. Lacking explicit awareness of program-level dependencies, Praxis (Raw Code) cannot effectively focus its reasoning on code regions relevant to the incident. Instead, it “fixated” on ostensibly faulty regions—code that appears erroneous on the surface (e.g., manually thrown exceptions)—that are not pertinent to the observed failure. With code context exceeding 200k tokens at times, Praxis (Raw Code) exhibited needle in a haystack [gola_multi_needle_2024] and context rot [hong2025contextrot] phenomena, leading to degraded RCA accuracy. These findings confirm that PDG-guided diagnosis is essential: it enforces the alignment of observability data with concise, accurate code context structurally presented to the LLM, grounding the RCA in relevant code paths and code regions to yield the observed performance gains.
@@ VI-C Diagnosis Overheads (RQ3)
For this RQ, we evaluated the RCA overheads of SRE-Agent baselines, Praxis, Praxis (Obs. Ctx.), and Praxis (Raw Code) using mean time to completion (MTTC) in seconds and average token consumption (ATC) in thousands of tokens, measured across all scenarios using the best-performing LLMs.
Evaluation setup. To account for differences in RCA reasoning accuracy, we normalized both MTTC and ATC by RCR Pass@1% to assess diagnosis efficiency: $\text{{Normalized metric}}=\frac{\text{{MTTC or ATC}}}{\text{RCR Pass@1\%}}$. The resulting normalized MTTC (i.e., mean time to diagnosis or MTTD) and normalized ATC (Effective ATC) reflect time and token costs per successful diagnosis. Here, MTTC is the average time to complete a diagnosis, regardless of correctness. MTTD is the normalized MTTC and represents the expected time to reach a correct diagnosis. Thus, MTTD jointly captures duration and accuracy: an agent that completes diagnosis faster but is less accurate may still require more time, in expectation, to reach a correct diagnosis, resulting in an MTTD longer than that of a more accurate agent. The same interpretation applies to ATC and Effective ATC.
Table˜V shows that although SRE-Agent has lower raw MTTC and ATC than Praxis, their MTTDs, which include LLM backend response times, are comparable. However, ReAct consumed $5.1\times$ more tokens per successful diagnosis, indicating much lower token efficiency. When we compare Praxis with its ablated variants (Praxis (Raw Code)) and Praxis (Obs. Ctx.)), construction of program context via LLM-driven PDG traversal reduced MTTD and effective ATC by up to $85.9\%$ and $83.5\%$, respectively, highlighting their significant contributions to efficient RCA. Notably, LLM-driven PDG traversal reduced ATC by $6.1\times$ compared to providing raw code as plain text in the prompt, as in [rcacp_chen2024, li2025coca_icse]. Praxis’s MTTD of 1,474 seconds is substantially shorter than that of manual RCA, which can take several hours for incidents with similar root causes.
@@
  • Metrics: (1) MTTC = mean time to completion in seconds; (2) ATC = average token consumptions in tokens; (3) MTTD = mean time to diagnosis; (4) Eff. ATC = effective ATC. Best-performing variant is shown in bold; second-best is underlined. Lower is better.
VI-D Discussions and Future Work
@@
Scenarios with weak observability. Praxis’s effectiveness is contingent on the availability and quality of observability data to anchor its PDG traversal. Sparse or ambiguous traces, logs, metrics, or events can limit Praxis’s ability to accurately align runtime error signals with corresponding dependency paths to traverse. When a fault offers no direct or clear observable signal at the application layer to anchor RCA decision-making, Praxis tends to falsely attribute the incident to fault-free code paths or configurations, leading to false positives. While collecting exhaustive observability data could mitigate this, it would incur significant production overheads. We explicitly identify the fault categories most vulnerable to this limitation in our scope discussion below.
Graph representation. Praxis’s effectiveness relies on the correctness and completeness of the dependency graphs. Dependency graphs that are incorrect or incomplete (e.g., an outdated PDG, or a missing microservice in SDG) hinder Praxis’s ability to diagnose a root cause. Because Praxis diagnoses faults by traversing dependency edges anchored on observability signals, missing edges break the structural path from symptom to root cause: observability may reveal the symptomatic component, but without the corresponding edge, the agent cannot reach the true root-cause node. For isolated nodes or a complete absence of dependency structures (e.g., infrastructure/resource dependencies), traversal terminates prematurely at the upstream symptom, reducing RCA accuracy. These implications explain Praxis’s practical scope and limitations in its present form, described in the next paragraph.
Praxis’s scope and limitations. Our evaluation (Table˜IV) shows that Praxis is effective at diagnosing (1) code-related incidents that fall under the fault models data schema mismatches, logical bugs, failure to handle external dependencies, and resource/address/label mismatches, and (2) misconfiguration-related incidents caused by misconfigured variables/constants in code or in ConfigMaps (see §VI-BTable˜IV). Collectively, these fault models represent the vast majority of code-related and misconfiguration incidents observed in production [ghosh_2022_socc, hotos_19, gunawi2014cloudbug]. Praxis works well here as these faults emit observability signals and either are directly visible in code or can be more accurately pinpointed via code examination (e.g., a misconfiguration altering execution paths).
As discussed, Praxis’s effectiveness depends on the availability and quality of observability signals and the completeness of its dependency graphs. Weak observability signals (e.g., missing traces or ambiguous logs) and incomplete or incorrect graphs can hinder diagnosis. This explains why Praxis and its variants achieve near-zero accuracy on deployment-manifest and resource/infrastructure-related incidents: those incidents lack direct application-layer observability and their dependency structures are not captured in Praxis’s graphs. We exclude these incidents from Table˜IV because all ablated variants behave and perform identically, so including them would not change the ablation comparison.
For the same reason, we consider concurrency/timing bugs out of scope: Praxis in its current form lacks dynamic runtime analysis to observe microservices’ internal runtime state and does not encode timing-dependency structures between microservices. These limitations are not inherent to the graph-based approach, and Praxis can be extended to address them, as discussed in the “Future work” section. Finally, while Praxis may generalize to fault types beyond those evaluated, its effectiveness on such faults remains unverified.
Access to source code. Praxis’s gains from PDG traversal are naturally limited when microservice code is unavailable for graph construction or code access is substantially limited (e.g., extensive reliance on proprietary third-party libraries), in which case Praxis must rely primarily on observability signals for diagnosis. In practice, partial code access is common (and matches our evaluation): Devs/SREs can access code developed in-house but not proprietary third-party binaries. Even with partial code access, Praxis can exploit clues in the available code captured by the PDG (e.g., call sites and configuration-controlled branches) to localize the root cause to external dependencies, thereby achieving higher RCA effectiveness than observability-only variants, see Figure˜1 and §VI-B for examples.
@@
Static code analysis. An inherent limitation of static code analysis is its inability to accurately capture dynamic bindings (e.g., Java reflection), which can leave unlinked gaps in the PDG. Although Praxis cannot eliminate this limitation, its hierarchical hammock-block traversal alleviates it by leveraging higher-level structural relations (e.g., class-level) to potentially bypass missing links at lower levels. Conversely, conditioning too strongly on static code context risks unintended bias, where Praxis may mistakenly attribute faults to code blocks when the incident actually stems from external factors (e.g., deployment or resources). This risk can be mitigated by incorporating runtime instrumentation to confirm whether implicated code paths were indeed executed at the time of the incident.
PDG construction overheads and continuous deployment. Regenerating all PDGs in our case took 116.42 seconds. Praxis builds PDGs by using a syntactic parser (e.g., Tree-sitter) to extract hammock blocks, and running CLDK-provided static analysis (e.g., WALA/RTA for Java and CodeQL/Scalpel for Python) [rahul2024cldk]. In practice, construction time is dominated by the static-analysis backend, and lightweight backends (e.g., the RTA [bacon1996rta] backend used by CLDK) scale near-linearly with microservice program size [grove_2001_rta]. Subsequent PDG construction takes $O(N+E)$, where $N$ is the number of hammock blocks and $E$ is the number of dependency edges extracted. Going forward, Praxis will be integrated with a CI/CD [^6] CI/CD stands for continuous integration and continuous deployment[redhat_cicd_2025]. pipeline. Praxis can avoid fully regenerating the PDGs because (1) PDGs can be constructed and parallelized per microservice, and (2) each PDG update can be incremental [hammer2015increment, souter2001incre], recomputing only nodes and transitive dependence regions affected by code changes. Moreover, caching versioned PDGs by commit SHA enables fast reuse and comparison.
Extensibility to other languages and platforms. Praxis is extensible beyond Python/Java and Kubernetes. SDG construction requires discovery of the runtime inter-service topology and dependencies, capabilities already supported by observability stacks (e.g., OpenTelemetry, Datadog) across platforms (e.g., OpenShift, ECS). PDG generation requires a syntactic parser for extracting hammock blocks and a static-analysis backend to derive control/data-flow edges; thus, specific language support is primarily limited by the availability of such toolchains. Because Praxis constructs PDG using Tree-sitter [tree_sitter_zenodo_2025] and CLDK [rahul2024cldk], which provide broad multi-language support, it can be readily extended to additional languages, including Go, Rust, TypeScript, and C/C++.
Future work. Future iterations of Praxis will be extended to diagnose resource-, infrastructure-, and deployment-related incidents by augmenting the current dependency graphs with resource/infrastructure and deployment-manifest graphs that explicitly capture those dependencies. Furthermore, Praxis will incorporate dynamic program analysis to address static-analysis gaps and enable the diagnosis of concurrency bugs. By leveraging lightweight, strategic runtime instrumentation (e.g., at database calls and gRPC handlers) alongside dynamic tracing with sampling [Liu2017DCatch, lu2018cloudraid, yuan2020effcon], Praxis can capture dynamic bindings and gain crucial visibility into internal microservice states. This runtime context will help reduce false positives for resource, infrastructure, and deployment faults, while simultaneously bringing concurrency bugs into scope. Ultimately, these techniques will ground the static graph traversal in verifiable runtime behavior, enabling more comprehensive incident coverage and precise RCA.
@@ VII Related Work
Agentic approach for cloud incident RCA. Recent agentic RCA approaches [roy2024exploring_rca_agents, jha2025itbench, xu2025openrca, chen2025aiopslab, chen2025stratus] adopted the ReAct paradigm [yao2023react] to analyze observability data but lack structured reasoning workflow and program context, limiting their effectiveness on software and misconfiguration faults. Prior approaches that incorporate program or code context [wang2024rcagent, li2025coca_icse] treat code as plain text in prompts rather than explicitly exploiting its inherent structure, as they do not perform PDG-guided reasoning. Praxis advances beyond those efforts by performing LLM-driven reasoning over service dependency and program dependence graphs, unifying incident, observability, cloud, and program context for scalable, end-to-end RCA.
@@ VIII Conclusion
This paper proposes Praxis, an agentic approach that reasons over microservice and program dependency graphs for a comprehensive, in-depth root-cause analysis. Our evaluation of Praxis on 30 scenarios that span software, configuration, deployment, and resource failures demonstrates that Praxis achieves up to $6.3\times$ better root-cause reasoning accuracy, $3.4\times$ higher entity identification accuracy, and a $5.3\times$ reduction in token consumption compared to state-of-the-art agentic baselines.
@@ Acknowledgment
We thank the reviewers, and R. Arora, Bhavya, N. Zheutlin, P. T. Isaza, J. Ahn, A. Paradkar, L. Shwartz, R. K. Kottapalli, A. D. Angelis, A. Patke, P. Cao, Z. Zheng, and J. Applequist for technical input and feedback, and H. C. Fairow, S. Weick, K. Atchley, Y. Deng, R. Pavuluri, M. Vukovic, X. Liu, D. Sow, and N. Fuller for administrative support. This work is supported by the IBM-Illinois Discovery Accelerator Institute (IIDAI) and NSF grant 2530738. Any opinions expressed here are those of the authors and do not necessarily reflect the views of IBM or NSF.